### v1

# SegEarth-OV3: Exploring SAM 3 for Open-Vocabulary Semantic Segmentation in Remote Sensing Images

Abstract
Most existing methods for training-free Open-Vocabulary Semantic Segmentation (OVSS) are based on CLIP. While these approaches have made progress, they often face challenges in precise localization or require complex pipelines to combine separate modules, especially in remote sensing scenarios where numerous dense and small targets are present. Recently, Segment Anything Model 3 (SAM 3) was proposed, unifying segmentation and recognition in a promptable framework. In this paper, we present a preliminary exploration of applying SAM 3 to the remote sensing OVSS task without any training. First, we implement a mask fusion strategy that combines the outputs from SAM 3’s semantic segmentation head and the Transformer decoder (instance head). This allows us to leverage the strengths of both heads for better land coverage. Second, we utilize the presence score from the presence head to filter out categories that do not exist in the scene, reducing false positives caused by the vast vocabulary sizes and patch-level processing in geospatial scenes. We evaluate our method on extensive remote sensing datasets. Experiments show that this simple adaptation achieves promising performance, demonstrating the potential of SAM 3 for remote sensing OVSS. Our code is released at https://github.com/earth-insights/SegEarth-OV-3.
1 Introduction
Semantic segmentation is a fundamental task in remote sensing analysis, enabling dense, pixel-level understanding of Earth observation scenes. Traditionally, segmentation models were restricted to a closed set of predefined categories, limiting their applicability in dynamic, open-world scenarios where visual concepts are virtually infinite. To overcome this limitation, Open-Vocabulary Semantic Segmentation (OVSS) has emerged as a critical research direction [40]. By leveraging the rich semantic knowledge embedded in pre-trained Vision-Language Models (VLMs), remote sensing OVSS aims to segment and recognize image regions based on arbitrary text descriptions, effectively generalizing to categories unseen during training. This capability is essential for diverse applications, e.g., urban planning [37] and disaster monitoring [57], where the model must handle a vast vocabulary [40, 39, 38].
Currently, the dominant paradigm for training-free OVSS relies heavily on VLMs, particularly CLIP [48]. Early works, such as MaskCLIP [72] and SCLIP [55], attempt to extract dense features directly from the CLIP image encoder. However, CLIP is pre-trained for image-level classification, and adapting its patch-level features for pixel-level localization often results in coarse boundaries. To address this, subsequent research has focused on integrating auxiliary Visual Foundation Models (VFMs). For instance, ProxyCLIP [34] and CorrCLIP [70] utilize structural guidance from DINO [47] and SAM [32] to refine CLIP’s attention maps, while SegEarth-OV [40, 39] builds an upsampler to reconstruct high-resolution features. Although these methods improve boundary quality, they rely on complex pipeline and feature alignment to bridge the gap between different representations. Moreover, they lack simultaneous semantic and instance segmentation capabilities, limiting their utility in complex geospatial analysis.
Recently, the Segment Anything Model 3 (SAM 3) [10] was introduced. Unlike CLIP-based paradigm, SAM 3 is a unified model that supports promptable concept segmentation. It builds upon the DETR [9] and MaskFormer [16, 17] architectures, employing a query-based Transformer design. Crucially, SAM 3 utilizes a decoupled architecture in which a presence head is specifically designed to predict the probability that the prompted concept exists in the image. Meanwhile, a Transformer decoder and a semantic segmentation head generate precise masks for discrete instances and continuous semantic regions, respectively. Although SAM 3 demonstrates impressive zero-shot capabilities on some natural image semantic segmentation benchmarks, remote sensing images present distinct challenges, e.g., the intricate coexistence of dense small objects and vast amorphous backgrounds. Therefore, a tailored exploration to adapt SAM 3 for geospatial scenarios remains valuable.
In this paper, we present a preliminary exploration of adapting SAM 3 for the remote sensing OVSS task without additional training. We investigate whether SAM 3’s unified architecture can offer a stronger, yet simpler baseline than complex CLIP-ensemble methods for Earth observation. Our proposed method, namely SegEarth-OV3, consists of two simple strategies tailored to SAM 3’s design:
(1) Dual-Head Mask Fusion: Remote sensing images exhibit a distinct duality: amorphous “stuff” (e.g., road, bareland) requiring pixel-wise semantic continuity, and countable “things” (e.g., buildings, vehicles) demanding instance-level boundary precision. We identify that SAM 3’s decoupled architecture naturally aligns with this duality. We propose to assign the semantic head to maintain land-cover completeness and the Transformer decoder to capture fine-grained instance details, thereby unifying these complementary strengths to ensure robust segmentation across diverse geospatial targets.
(2) Presence-Guided Filtering: In remote sensing OVSS, a complete vocabulary list might include global land cover types, but a single image patch covers a minute geographical area (e.g., hundreds of meters). This results in a high category sparsity, where the vast majority of queried concepts are physically absent in the local view. We leverage the presence score to address this global-local discrepancy, explicitly suppressing irrelevant categories to eliminate false positives caused by the vast vocabulary against limited visual content [53, 14].
We evaluate our approach on 17 remote sensing semantic segmentation benchmarks and some general segmentation benchmarks. Our results demonstrate the strong capability of SAM 3 for remote sensing OVSS, which is further enhanced by our proposed improvements.
2 Related Work
denotes the element-wise maximum operation, and
denotes multiplication.2.1 Training-based OVSS
Training-based OVSS methods fine-tune pre-trained VLMs on annotated datasets, typically adopting either a mask classification or a dense feature adaptation paradigm. Mask classification methods, including OpenSeg [24], OVSeg [41], ZegFormer [21], and MasQCLIP [66], etc., leverage generated class-agnostic masks for subsequent CLIP classification. In contrast, dense feature adaptation methods, ranging from early pixel-alignment works like LSeg [35] to advanced adapter-based models like SAN [65], SED [62] and CAT-Seg [18], refine CLIP’s feature maps directly for dense prediction via side networks or cost aggregation. Other methods such as SegCLIP [43] and GroupViT [64] explore weakly-supervised grouping from image-text pairs. Despite achieving strong performance, these methods require computationally expensive training on large-scale datasets, limiting their flexibility compared to training-free alternatives.
2.2 Training-free OVSS
To avoid training costs, training-free methods directly adapt pre-trained VLMs for dense prediction. Pure CLIP-based methods such as SCLIP [55] and ClearCLIP [33] modify the internal mechanisms of CLIP, e.g., removing pooling layers or refining self-attention maps, to extract dense features [72, 5, 26, 29]. However, due to CLIP’s image-level pre-training objective, these methods often suffer from coarse localization. To address this, the VFM-assisted methods [34, 70, 51, 31, 3] integrate auxiliary models like DINO [11, 47] or SAM [32, 49] to provide structural guidance. By utilizing attention maps or object proposals from these foundation models, they achieve clearer boundaries. However, these methods operate as disjointed pipelines, relying on separate, heavy models for segmentation and recognition, which leads to significant system complexity.
2.3 Remote Sensing OVSS
Extending OVSS to remote sensing images faces unique challenges, such as extreme scale variations and arbitrary orientations, which often degrade the performance of methods designed for natural images. To mitigate the coarse localization of CLIP, recent methods like SegEarth-OV [40] and GSNet [69] employ feature upsampling modules or dual-stream architectures to incorporate domain-specific priors, while SkySense-O [74] advances this by pre-training vision-centric VLMs on large-scale remote sensing data. Addressing the specific geometric complexities of aerial views, Cao et al. [8] introduce rotation-aggregative modules to handle orientation diversity. Furthermore, AerOSeg [22] and SCORE [27] leverage SAM features for structural guidance and regional context, while RemoteSAM [68] and InstructSAM [71] utilize SAM-based pipelines to unify segmentation with broader interpretation tasks. However, these methods typically rely on complex multi-stage pipelines or require expensive domain-specific training. In contrast, we explore SAM 3 as a unified, training-free framework to simplify remote sensing OVSS.
3 Methods
3.1 Preliminaries
We adopt SAM 3 as our foundational architecture. Unlike standard semantic segmentors that map an image directly to a label map , SAM 3 operates as a prompt-conditioned predictor. Given an image and a specific text prompt (e.g., a category name “building”), the model predicts the probability that a pixel or region belongs to the concept defined by . Architecturally, SAM 3 consists of a vision encoder and a text encoder that extract image embeddings and text embeddings, respectively. These are processed by a Fusion Encoder, which outputs prompt-conditioned image features . Based on , the model utilizes three decoupled heads to generate predictions:
Presence Head: Predicts a scalar score , indicating the global probability that the concept exists in the image.
Semantic Segmentation Head: A dense prediction module (FCN-style) that maps to a semantic probability map .
Transformer Decoder (Instance Head): A query-based module that outputs a set of instance predictions . Here, is the probability map for the -th object query, and is its associated confidence score.
For OVSS, a naive baseline is to rely solely on the instance predictions from the Transformer decoder [10], aggregating them to form a segmentation mask. However, in remote sensing, this baseline often overlooks amorphous regions and is prone to false positives when querying a large vocabulary. Motivated by this, we introduce some strategies to address these limitations.
3.2 Inference Pipeline
We propose a training-free inference strategy designed to tackle the specific challenges of OVSS in remote sensing scenarios, as shown in Figure 2. Our pipeline processes each category in the vocabulary sequentially and aggregates the results into a final segmentation map.
Instance Aggregation. Remote sensing scenes often contain dense clusters of small, identical objects (e.g., vehicles in a parking lot, ships in a harbor). Standard semantic segmentation methods often leads to boundary adhesion in dense clusters. To address this, we first leverage the Transformer decoder, which treats objects as discrete queries, effectively isolating individual instances even in crowded scenes. It generates a set of discrete instance predictions . We aggregate these sparse predictions into a single category-level map by taking the maximum weighted probability at each pixel:
This aggregation effectively consolidates individual object instances into a unified semantic layer, preserving the fine-grained localization capabilities of the instance head even in crowded scenes.
Dual-Head Mask Fusion. While the instance head excels at delineating countable objects (“things”), it may produce fragmented predictions for large-scale continuous and amorphous regions (“stuff”) like road or bareland, which are prevalent in remote sensing images, as shown in Figure 1. Conversely, the semantic head provides dense, global coverage but often blurs the boundaries of small targets or misses them entirely. To reconcile these complementary strengths, we fuse the aggregated instance map with the dense probability map from the semantic head using a max-fusion strategy:
This fusion ensures robust segmentation performance across diverse categories, capturing both the distinct boundaries of small instances and the completeness of large-scale amorphous regions.
Presence-Guided Filtering. A critical challenge in OVSS arises from querying a massive global vocabulary against a localized image patch. Although the vocabulary might enumerate a comprehensive list of global land-cover types (e.g., various biomes, infrastructure), a single inference patch restricts the view to a small geographical extent (e.g., a few hundred meters). This results in a high category sparsity, where the valid targets in a given image are only a small subset of . The model becomes prone to hallucinating absent categories due to textural ambiguities common in geospatial data, e.g., confusing “barren land” with “sports field”. To mitigate this, we utilize the global presence score to explicitly suppress irrelevant categories. We apply a soft gating operation, , which reduces the weight of the probability maps of categories predicted to be absent. Finally, we assign each pixel to the category with the highest probability:
To handle ambiguous regions, pixels with a maximum probability below a threshold are assigned to the “background” category (if it exists). Note that for categories with multiple prompts, we select the one with the highest probabilities to ensure robustness.
4 Experiments
| Methods | OpenEarthMap | LoveDA | iSAID | Potsdam | Vaihingen | UAVidimg | UDD5 | VDD | Avg |
[…]
| SCAN [42] | - | 23.2 | 44.3 | 27.5 | 15.2 | 20.3 | 34.1 | 29.2 | - |
| SAN [65] | - | 25.3 | 49.6 | 37.3 | 39.2 | 23.5 | 37.2 | 35.8 | - |
| SED [62] | - | 24.6 | 51.2 | 29.4 | 39.0 | 21.3 | 35.7 | 32.5 | - |
| Cat-Seg [18] | - | 28.6 | 53.3 | 35.8 | 42.3 | 25.7 | 40.2 | 39.1 | - |
| OVRS [8] | - | 31.5 | 52.7 | 36.4 | 43.5 | 24.1 | 40.8 | 37.2 | - |
| GSNet [69] | - | 32.5 | 53.7 | 37.9 | 44.1 | 24.2 | 40.9 | 37.3 | - |
| RSKT-Seg [36] | - | 33.2 | 54.3 | 38.4 | 42.7 | 25.7 | 42.1 | 39.7 | - |
| SkySense-O [74] | 40.8 | 38.3 | 43.9 | 54.1 | 51.6 | - | - | - | - |
[…]
| CLIP [48] | 12.0 | 12.4 | 7.5 | 15.6 | 10.8 | 10.9 | 9.5 | 14.2 | 11.4 |
| MaskCLIP | 25.1 | 27.8 | 14.5 | 33.9 | 29.9 | 28.6 | 32.4 | 32.9 | 27.2 |
| SCLIP [55] | 29.3 | 30.4 | 16.1 | 39.6 | 35.9 | 31.4 | 38.7 | 37.9 | 31.1 |
| GEM [5] | 33.9 | 31.6 | 17.7 | 39.1 | 36.4 | 33.4 | 41.2 | 39.5 | 32.3 |
| ClearCLIP [33] | 31.0 | 32.4 | 18.2 | 42.0 | 36.2 | 36.2 | 41.8 | 39.3 | 33.4 |
| SegEarth-OV [40] | 40.3 | 36.9 | 21.7 | 48.5 | 40.0 | 42.5 | 50.6 | 45.3 | 39.2 |
| ProxyCLIP [34] | 38.9 | 34.3 | 21.8 | 49.0 | 47.5 | 35.8 | 40.8 | 47.8 | 39.5 |
| CorrCLIP [70] | 32.9 | 36.9 | 25.5 | 51.9 | 47.0 | 38.3 | 46.1 | 47.3 | 40.7 |
| SegEarth-OV3 | 42.9 | 47.4 | 27.6 | 57.8 | 60.8 | 54.7 | 71.7 | 64.5 | 53.4 |
| Oracle | 64.4 | 50.0 | 36.2 | 74.3 | 61.2 | 59.7 | 56.5 | 62.9 | 58.2 |
| Method | Building Extraction | Road Extraction | Flood Detection | ||||||
| WHUAerial | WHUSat.II | Inria | xBDpre | CHN6-CUG | DeepGlobe | Massachusetts | SpaceNet | WBS-SI | |
| CLIP [48] | 17.7 | 3.5 | 19.6 | 16.0 | 7.7 | 3.9 | 4.9 | 7.1 | 18.6 |
| MaskCLIP [72] | 29.8 | 14.0 | 33.4 | 29.2 | 28.1 | 13.2 | 10.6 | 20.8 | 39.8 |
| SCLIP [55] | 33.4 | 21.0 | 34.9 | 25.9 | 21.1 | 7.0 | 7.4 | 14.9 | 32.1 |
| GEM [5] | 24.4 | 13.6 | 28.5 | 20.8 | 13.4 | 4.7 | 5.1 | 11.9 | 39.5 |
| ClearCLIP [33] | 36.6 | 20.8 | 39.0 | 30.1 | 25.5 | 5.7 | 6.4 | 16.3 | 44.9 |
| SegEarth-OV [40] | 49.2 | 28.4 | 44.6 | 37.0 | 35.4 | 17.8 | 11.5 | 23.8 | 60.2 |
| SegEarth-OV3 | 86.9 | 44.2 | 72.4 | 64.3 | 49.6 | 39.3 | 27.7 | 35.6 | 75.6 |
4.1 Setup
Datasets. Following SegEarth-OV [40], we evaluate our method on 17 remote sensing datasets, covering diverse scenes, resolutions, and tasks.
We further test on nine binary extraction datasets focusing on critical geospatial objects: building extraction (WHUAerial [28], WHUSat.II [28], Inria [45], and xBD [25]), road extraction (CHN6-CUG [73], DeepGlobe [20], Massachusetts [46], and SpaceNet [54]), and flood detection (WBS-SI222https://www.kaggle.com/datasets/shirshmall/water-body-segmentation-in-satellite-images).
Additionally, we report results on general scene datasets (Pascal VOC20 [23], COCO Stuff [6], and Cityscapes [19]) to demonstrate universality. The evaluation metric is mIoU for multi-class segmentation and IoU of the foreground class for binary extraction tasks.
Implementation Details. We use the official SAM 3 model equipped with the Perception Encoder-Large+ (PE-L+) [4] backbone. Input images are resized to 10081008. Text prompts are derived directly from category names (e.g., “building”, “road”), and some categories are augmented with synonyms333The prompt setting is manually curated based on dataset features and has not been systematically explored in this version.. The background threshold and the initial confidence threshold for the Transformer decoder are manually tuned for each dataset to achieve roughly optimal performance. No test-time augmentation and extra post-processing method [1] is used.
4.2 Main Results
Semantic segmentation. We report the quantitative comparison on eight remote sensing semantic segmentation benchmarks in Table 4. Our method, SegEarth-OV3, achieves a new state-of-the-art, demonstrating a substantial performance leap over existing methods. Specifically, it achieves an average mIoU of 53.4%, surpassing the best previous training-free method, CorrCLIP (40.7% mIoU), by a remarkable margin of +12.7% mIoU. This performance also consistently outperformed training-based OVSS methods fine-tuned on remote sensing data, e.g., RSKT-Seg (39.7% mIoU) and Cat-Seg (39.1% mIoU). A notable observation is that our zero-shot method outperforms the fully supervised Oracle on some datasets. On UDD5 and VDD datasets, SegEarth-OV3 achieves 71.7% mIoU and 64.5% mIoU respectively, exceeding the Oracle baselines (56.5% and 62.9% mIoU). This surprising result suggests that the rich semantic knowledge and robust segmentation capabilities inherent in the SAM 3 foundation model can, in certain scenarios, exceed the generalization ability of models trained on domain-specific data. In Figure 4, we visualize the results on a large-scale remote sensing image.
| Method | LoveDA | Uavid | xBDpre | CHN6-CUG |
| Instance Only | 32.2 | 50.4 | 61.4 | 38.4 |
| Semantic Only | 35.4 | 47.1 | 44.9 | 39.5 |
| SegEarth-OV3 | 47.4 | 54.7 | 64.3 | 49.6 |
| Method | Size | VOC20 | Stuff | City |
[…]
| TCL [12] | ViT-B/16 | 83.2 | 22.4 | 24.0 |
| CLIP-DINOiser [60] | 80.9 | 24.6 | 31.7 | |
| CoDe [59] | - | 23.9 | 28.9 | |
| CAT-Seg [18] | 94.6 | - | - | |
[…]
| CLIP =[48] | ViT-B/16 | 41.9 | 4.4 | 5.0 |
| MaskCLIP [72] | 74.9 | 16.4 | 12.6 | |
| ClearCLIP [33] | 80.9 | 23.9 | 30.0 | |
| SCLIP [55] | 80.4 | 22.4 | 32.2 | |
| ProxyCLIP [34] | 80.3 | 26.5 | 38.1 | |
| LaVG [30] | 82.5 | 23.2 | 26.2 | |
| CLIPtrase [50] | 81.2 | 24.1 | - | |
| NACLIP [26] | 83.0 | 25.7 | 38.3 | |
| Trident [51] | 84.5 | 28.3 | 42.9 | |
| ResCLIP [67] | 86.0 | 24.7 | 35.9 | |
| SC-CLIP [2] | 84.3 | 26.6 | 41.0 | |
| CLIPer [52] | 85.2 | 27.5 | - | |
| CASS [31] | 87.8 | 26.7 | 39.4 | |
[…]
| FreeDA [3] | ViT-L/14 | 87.9 | 28.8 | 36.7 |
| CaR [53] | 91.4 | - | - | |
| ProxyCLIP [34] | 83.2 | 25.6 | 40.1 | |
| ResCLIP [67] | 85.5 | 23.4 | 33.7 | |
| SC-CLIP [2] | 88.3 | 26.9 | 41.3 | |
| CLIPer [52] | 90.0 | 28.7 | - | |
[…]
| ProxyCLIP [34] | ViT-H/14 | 83.3 | 26.8 | 42.0 |
| Trident [51] | 88.7 | 28.6 | 47.6 | |
[…]
| SegEarth-OV3 | PE-L+/14 | 96.8 | 42.8 | 69.7 |
Single-class extraction. We further evaluate the single-class land-cover extraction capability of our method on nine benchmarks focusing on buildings, roads, and floods (Table 2). In building extraction, our method achieves unprecedented gains, reaching 86.9% IoU on WHUAerial and 72.4% IoU on Inria, outperforming the previous state-of-the-art SegEarth-OV by massive margins of +37.7% and +27.8% respectively. Similarly, for road extraction, it consistently surpasses previous methods, achieving 49.6% IoU on CHN6-CUG. In the flood detection task, SegEarth-OV3 attains 75.6% IoU, marking a +15.4% improvement. These consistent and dramatic improvements across diverse geospatial targets, emphasize the robustness of SegEarth-OV3 and the generalization capability of SAM 3.
4.3 Ablation Studies
We investigate the contribution of the dual-head fusion strategy by evaluating the performance of each head individually on four representative datasets, as listed in Table 3. Using only Transformer decoder predictions yields strong results on object-centric tasks like building extraction, but underperforms on complex scenes, due to struggles with amorphous regions. Conversely, the “Semantic Only” baseline struggles with precise object boundaries, achieving only 44.9% IoU on xBD. By fusing both heads, SegEarth-OV3 achieves significant improvements. On LoveDA and CHN6-CUG, SegEarth-OV3 reaches 47.4% mIoU and 49.6% IoU, surpassing the best single-head baselines by +12.0% and +10.1%. Even on the instance-heavy xBD dataset, fusion further boosts performance by 2.9%. These results validate our hypothesis that the semantic and instance heads provide complementary information, i.e., global coverage and fine-grained precision, and their combination is essential for remote sensing OVSS.
4.4 Results on General Scene Datasets
We benchmark SegEarth-OV3 on three standard general scene datasets: Pascal VOC20, COCO Stuff, and Cityscapes (Table 4.2). For Pascal VOC20 and Cityscapes, we refine category names to better align with visual concepts; for instance, the “terrain” class in Cityscapes is expanded to “grass, horizontal vegetation, soil, sand” based on official definitions444https://www.cityscapes-dataset.com/dataset-overview. SegEarth-OV3 achieves dominance across all benchmarks. On Pascal VOC20, it achieves 96.8% mIoU, surpassing both the best training-free method CorrCLIP (91.8% mIoU) and the training-based CAT-Seg (94.6% mIoU). On COCO Stuff, it reaches 42.8% mIoU, significantly exceeding CorrCLIP by +8.8%. The most significant improvement is observed on Cityscapes, where our method achieves 69.7% mIoU, representing an increase of 18.6% mIoU over the previous best result. These results further emphasize the powerful capabilities of SAM 3 and the effectiveness of SegEarth-OV3.
5 Conclusion
This paper explores the potential of SAM 3 for remote sensing OVSS. Building upon the powerful zero-shot capabilities of the SAM 3 foundation model, we introduce an inference strategy to adapt its prompt-based architecture for remote sensing OVSS. By fusing semantic and instance outputs and leveraging presence-guided filtering, we effectively address the unique challenges of remote sensing. Our results on 20 diverse datasets show that SegEarth-OV3 not only sets a new training-free state-of-the-art but also outperforms supervised baselines in specific cases. These results highlight the significant potential of SAM 3 for specific domain tasks and demonstrate that, through appropriate adaptation strategies, training-free paradigms can serve as a powerful alternative to traditional supervised learning.